Status Report

ISS On-Orbit Status 20 Aug 2002

By SpaceRef Editor
August 20, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below.

Onboard ISS the night has turned to day: Crew day again began yesterday afternoon at 4:00 pm EDT, ending 7:30 am this morning. Wake-up tonight is at 4:00 pm.

The two cosmonauts spent several hours with systems servicing tasks. CDR Korzun completed periodic preventive maintenance of the Service Module (SM) ventilation system, specifically checking and cleaning fans of Group A. FE-2 Treschev worked in the Docking Compartment (DC-1) "Pirs", first cleaning its air ducts (VD-1, VD-2), then replacing two spent dust collector filters (PF1, PF2). Later, working jointly, they also exchanged four SM dust filters (PF 1-4).

FE-1 Whitson completed the periodic calibration of the RED (resistive exercise device) canisters, partly assisted by Korzun, who then logged the data. The RED will reach the end of its certified life in 1-2 weeks. A set of spare Flexpak canisters is on board, and ground replacements will be manifested later.

Peggy Whitson also performed the more elaborate monthly TVIS treadmill inspection and maintenance. [Analysis of photographs taken by the crew before and after removal of the shaft from the damaged roller has raised some concerns on the ground about a possible crack in the chassis. It appears to be only a surface flaw, but the crew was asked to further inspect the location to make sure.]

Valery Korzun and Sergey Treschev completed another session with the Russian-Japanese HDTV (high definition television) experiment that provides close-up facial images of the crewmembers during TVIS exercise or normal conversation for subsequent medical/psychological analysis. The session was taped on the HDTV camcorder and is also downlinked to TsUP in real-time TV on UHF (ultrahigh frequency) some of the time, when the station was over RGS (Russian ground sites) on Daily Orbit 15.

Peggy took care of the regular daily routine tasks of SOSh life support systems maintenance in the SM, checkup of autonomous Lab payloads (PCG-STES008 & ADVASC status monitor), and IMS (inventory management system) delta file preparation.

After a smoke detector (SD) in the Lab EXPRESS Rack 2 (ER2) yesterday indicated an obscuration reading without apparent cause, it was recovered overnight and the reading has stabilized. MCC-H has re-enabled the SD, returning ER2 to nominal payload ops.

MCC-H also finished the open voltage reading of the recharged EMU batteries and determined that they are good. After both batteries have now been properly discharged and recharged, work can continue on troubleshooting the discharge problem of the BSA/BCA (battery storage assembly/battery charger assembly) in the Airlock module.

Power drain procedure began again on the #2 battery of the P6 module’s 4B channel, to continue for about three days. This is a periodic maintenance procedure to reduce the pressure differential between the battery cells to an acceptable level.

At 5:00 am EDT, Valery and Peggy engaged in the scheduled interactive audio/video interview with the NBC TV, which was taped and then played back later this morning on the Today Show

As per Moscow’s report to the MMT today, the 1h 43 min delay in EVA-7 egress on 8/16 was caused by CDR Korzun’s inadvertently skipping a crucial step in the Orlan suit preparation sequence for both suits, viz., opening the control valve of the backpacks’ primary O2 supply. [Procedural steps are being taken to safeguard against a repeat of the mistake (for example: the customary written procedures for suit preparation will henceforth be accompanied by a checklist, and each suit will be prepared sequentially, not both in parallel. Also, for EVA-8, Sergey Treschev will double-check after Korzun).]

To accommodate EVA-7 tasks not performed, EVA-8, the second spacewalk from DC-1, has been rescheduled from 8/23 to 8/26 (Monday). Any later date will aggravate problems with RGS and Russian satellite communications compatibility (between RGS and MCC-M). Hatch opening will be at about 1:00 am EDT, and duration about 6 h. [In preparing for the modified EVA at the new date, TsUP must re-negotiate with the Russian military for comsat channels. For 8/26, one channel has already been secured, which, along with US S-band, would suffice for the EVA. An extra channel, if obtained, would provide TsUP with additional command uplink. An EVA-8 Readiness Review is scheduled for the next MMT meeting on 8/23 (Friday).]

The Elektron oxygen generator continues having problems with nitrogen/air bubbles in its system, causing its pump to shut down after operating for half a day or so. After restart, the machine works OK again. Experience from Mir has shown that it takes the cavitating bubbles about two weeks to dissolve. The operational periods are currently trending up in length, but it remains unsure whether the Fluid Unit can be fully recovered. A new Fluids Unit on Progress 9P will resolve the issue if it persists.

MSG/SUBSA (Microgravity Science Glovebox/Solidification Using a Baffle in Sealed Ampoules) personnel has developed MSG work volume cleaning steps which are being reviewed by MSFC Safety specialists. An integrated procedure, being put together right now, may be ready for implementation tomorrow. [The quartz ampoule of sample SUBSA-07 cracked on 8/11 during cooldown in the MSG thermal chamber and subsequently broke, setting free its contents of Antimonide and its salt-like encapsulant. The crack was due to an unexpected temperature gradient in the Antimonide sample that caused it to solidify radially instead of axially (lengthwise), expanding in the process and bursting the ampoule.]

Yesterday, the MCC-H building suffered an electrical short during maintenance work, requiring reboot of several computers. Until a "clean" configuration was reestablished, ISS telemetry could not be received for 1 h 7 min.

Sincere congratulations were expressed today by ISS Mission Management Team (MMT) personnel for NASA’s Voyager team. Because on this day 25 years ago Voyager 2 began its epic voyage without end. [Currently at 68 AUs from Earth (6.3 Billion miles/10 billion km), its radio signals take over nine hours to reach us. (Voyager 1 was launched on 9/5/77 and, at 85 AUs, i.e. 8 billion mi./12.7 billion km, is the most distant human-made object.)]

Targets for the U.S. CEO (crew earth observations) program today were Angolan Biomass Burning (as the burning season progresses, fires should be spreading towards central and southeast Africa [Zambia and Zimbabwe, right of track] and reaching their peak in Angola [left of track]), Congo-Zimbabwe Biomass Burning (good conditions for observing fires and smoke remain. Oblique views left and right of track are best. Good view down the Lake Tanganyika rift right of track), Beijing, China (views left and right of track for this vast urban region), former typhoon Phanfone (former super typhoon that scared Tokyo in the last few days has veered back into the Pacific and was downgraded to a tropical storm by the time ISS flew over it), and Western Mediterranean Aerosol (smog in southern France, northern Italy [both left] and out over the Mediterranean [right of track]).

CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 4:00 am EDT):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
Elektron O2 generator is powered On (32-amp mode), on backup pump. Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5, i.e., 10-min. cycle time (vacuum pump failed). U.S. CDRA CO2 scrubber is on Override. BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.

Electrical Power Systems (EPS):
Both P6 channels fully operational. Beta Gimbal Assembly (BGA) 2B and BGA 4B in Autotrack mode (solar-tracking).
SM batteries: Battery #1 is off-line; all other batteries (7) are in "Complete Charge" mode.
FGB batteries: Battery #5 is off-line, battery #3 is in "Cycle" mode; all other batteries (4) are in "Complete Charge" mode.
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode (were in Discharge mode during EVA-7).

Thermal Control Systems:
Air conditioner SKV-1 is On. SKV-2 is Off.

Command & Data Handling Systems:
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1 is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is operational; PL-2 MDM is Off.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:
3 CMGs on-line.
State vector — US GPS (SIGI string 1)
Attitude — Russian segment
Angular rates — US RGA1 (rate gyro assembly 1)

Communications & Tracking Systems:
All Russian communications & tracking systems are nominal.
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is operating nominally.

SSRMS/Canadarm2 at MBS PDGF 1 (mobile base system/power & data grapple fixture 1), in EVA-7 viewing position, with Keep Alive power on both strings (based on MBS).
MBS: Keep Alive power on both strings.
RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning, 6:58 am EDT [= epoch]):
Mean altitude — 393.1 km
Apogee — 404.7 km
Perigee — 381.6 km
Period — 92.4 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0017065
Orbits per 24-hr. day — 15.58
Altitude decrease — 250 m (mean) in last 24 hours
Solar Beta Angle — -32.0 deg (magnitude increasing)
Revolutions since FGB/Zarya launch (Nov. ’98) — 21405
Current Flight Attitude — XPOP (x-axis perpendicular to orbit plane = "sun-fixed" [yaw: -0.6 deg, pitch: -5.7 deg., roll: 0 deg]).

For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.